# Qualitative Failures of Image Generation Models and Their Application in Detecting Deepfakes

### v1

@@ Abstract
The ability of image and video generation models to create photorealistic images has reached unprecedented heights, making it difficult to distinguish between real and fake images in many cases. However, despite this progress, a gap remains between the quality of generated images and those found in the real world. To address this, we have reviewed a vast body of literature from both academic publications and social media to identify qualitative shortcomings in image generation models, which we have classified into five categories. By understanding these failures, we can identify areas where these models need improvement, as well as develop strategies for detecting deep fakes. The prevalence of deep fakes in today’s society is a serious concern, and our findings can help mitigate their negative impact. In order to support research in this field, a collection of instances where models have failed is made available at here.
@@ 1 Introduction
Generated images, also known as synthetic images, are created by machine learning algorithms or other software programs, while real images are captured by cameras or other imaging devices. Generated images are not real-world representations of a scene or object, but rather computer-generated approximations. As such, they lack the authenticity and realism of real images.
@@
As the quality of generated images continues to improve, it is crucial to conduct more in-depth and precise analyses. Thus far, people have been amazed by the ability of synthesized images to approximate natural scenes. When Photoshop was introduced, significant efforts were made to identify manipulated images, and a similar approach is needed for generated images today. It would be beneficial to compile a set of indicators and other resources to aid in detecting deep fakes.
We present a collection of indicators that can be examined in a single image to determine whether it is genuine or manipulated. Overall, we offer five classes of these indicators including Human and Animal Body Parts, Geometry, Physics, Semantics and Logic, as well as Text, Noise, and Details, for both portraits and natural landscapes. The advantage of utilizing qualitative cues is that they are easily accessible and can be utilized by anyone, potentially serving as the initial step in detecting deep fakes.
@@ 2 Related work
2.1 Quantitative Approaches to Evaluate Generative Models
@@
Quantitative approaches have emerged as a vital tool to evaluate the performance of generative models. These methods rely on quantitative measures to assess how well a model is able to generate realistic data. One commonly used metric is the Inception Score [26], which evaluates the diversity and quality of generated images based on the classification accuracy of a pre-trained classifier. Another popular approach is the Fréchet Inception Distance [14], which uses feature statistics to compare the distribution of generated data with that of real data. Moreover, other metrics such as precision and recall [25] can be used to evaluate the quality of generated samples in specific domains such as vision, text and audio. These quantitative approaches provide a rigorous and objective way to measure the effectiveness of generative models, helping researchers to improve their models and develop more advanced generative techniques. Large-scale datasets have also been used in studies that focus on the qualitative evaluation of generated images (e.g.  [2]). To obtain a thorough analysis of quantitative metrics for evaluating generative models, please refer to the following references [4, 6, 28, 30].
@@ 2.2 Deepfake Detection Methods
Deepfake detection has become an essential area of research due to the increasing sophistication of deep learning algorithms that can generate highly realistic fake images, videos, and audio. As a result, numerous deepfake detection methods have been proposed in recent years, ranging from traditional image and video forensic techniques to advanced deep learning-based approaches. These methods can be broadly categorized into two groups: static and dynamic analysis.
@@ 3 Qualitative Failures of Image Generation Models
We compiled a list of qualitative failures by examining images from various sources including social media websites such as Twitter, LinkedIn, Discord, and Reddit [^1] A few of the images used in this work were obtained with the consent of a Reddit user named Kronzky., as well as images from the DiffusionDB dataset [29]. These images have been generated by notable generative models such as DALL-E 2, Midjourney, StableDiffusion, and Bing Image Creator. Additionally, we analyzed images from websites such as thisxdoesnotexist.com, whichfaceisreal.com, the Adobe Stock library, and openart.ai. We made sure that the text prompts used to generate images were not intentionally seeking peculiar images. Finally, we manually reviewed the images and filtered out the ones without problems.
@@ 3.1 Human and Animal Body Parts
Figure 2: Examples of poorly generated faces.
Image Background. When creating deepfakes, issues with the background of the images may arise, particularly in cases where the face is in focus while the surrounding clues are incorrect. The neural network used to generate the images focuses mainly on the face and may not pay as much attention to the surrounding details. This can lead to strange companions or chaotic forms in the background. Additionally, the objects or people next to the primary person in the image may appear unnatural or “mutant”. Figure 3 displays several instances of failures as examples.
@@
Figure 4: Here are some instances of eyes that were generated poorly. The eye in the bottom right corner is an actual photograph of a patient who has an irregularly shaped pupil. You can refer to this link for more details. This case represents a unique manifestation of a condition known as “cat’s eye Adie-like pupil,” which is considered a warning sign for ICE syndrome.
@@ 3.2 Geometry
Deepfakes may exhibit anomalous or atypical image geometry, with objects appearing to be of an unusual shape or size, in comparison to their expected proportions.
Straight Lines and Edges. AI-generated images may lack the straight lines, seams, and connections found in real-world objects, resulting in wavy, misaligned, and jumpy renderings (e.g. in tiles). Deepfakes can also exhibit inconsistent or unnatural image edges, which refer to the boundaries between different parts of the image. Further, surfaces, which are typically straight, may look somewhat uneven in generated images. Some samples failures are shown in Figure 14.
@@
Perspective. Models lack the ability to understand the 3D world, which results in physically impossible situations when objects cross different planes in a scene. These errors are difficult to detect as our brain often auto-corrects them, requiring a conscious investigation of each angle of the object to identify inconsistencies. Deepfakes can display an unnatural or distorted perspective, where a person’s body appears stretched or compressed unrealistically. Deepfakes may also have inconsistent or unrealistic camera angles, where a person’s face appears to be viewed from an impossible angle or perspective. Some example failures are shown in Figure 15.
@@
Other Geometry. Deepfakes exhibit various geometrical anomalies that may reveal their artificiality. For instance, their depth cues can be inconsistent or unnatural, causing the foreground or background to seem blurry or devoid of detail. Moreover, deepfakes often lack parallax, which is the apparent displacement of objects when viewed from different perspectives, resulting in a flat or two-dimensional appearance. Additionally, incorrect or inconsistent motion blur may suggest that certain parts of the image have been manipulated. The absence of occlusion, i.e., the overlapping of objects in the scene, is another telltale sign of deepfakes, as it can make the image look flat or unrealistic. Lastly, deepfakes may display improper image alignment, with objects seeming misaligned or out of place.
@@ 3.3 Physics
Reflection. An effective technique for detecting generated images is to examine the lighting and how it interacts with the elements within the image, and how it causes reflections and shadows. Deepfakes can exhibit artificial reflections that are inconsistent with the natural lighting and environment, such as those in glasses, mirrors, or pupils. The root cause of this issue is that deep generative models lack a proper understanding of reflections. While these models may recognize that an image contains a reflection and typically involves two people (one facing the camera and the other with their back turned), they do not comprehend that the two individuals are, in fact, the same person. Deepfakes may display other lighting effects that do not match real-world environments, such as lens flares, lens distortion, chromatic aberration and unnatural specular highlights. These effects are frequently observed in genuine photographs due to the physical properties of camera lenses and the way light is refracted through them. Some example failures are shown in Figure 19.
@@
Shadow. Deepfakes might not include shadows, which are typically found in real images and contribute to the impression of depth and authenticity. It is important to observe objects without shadows and those with highlights that appear to originate from a different direction than the rest of the image. Additionally, if the photo was taken outdoors in natural light during the afternoon, the setting sun will produce longer shadows than it would at midday, which can be easily identified by scrutinizing the shadow’s length. However, this method may not be as precise in artificial lighting conditions. Finally, if there are multiple objects or people within the scene, their shadows should be consistent with each other. Some generated images with inconsistent shadows are shown in 20.
@@ 3.4 Semantics and Logic
Figure 22: Samples of spatial reasoning from [20]. Images are generated by DALL-E 2 for the following text prompts for columns from left to right: “a red basketball with flowers on it, in front of blue one with a similar pattern”, “a red ball on top of a blue pyramid with the pyramid behind a car that is above a toaster”, “a pear cut into seven pieces arranged in a ring, “In late afternoon in January in New England, a man stands in the shadow of a maple tree”, and “An old man is talking to his parents”.
Context and Scene Composition. Deepfakes can be detected through various inconsistencies such as the background or surroundings not matching the real-world environment, missing contextual details, unnatural object placement, and inconsistent image composition. These irregularities may include inconsistencies in order of objects, missing objects or features, objects appearing in the wrong location or orientation, or unnatural arrangement and placement of objects in the image. Please see Figure 23.
@@
Other Semantics. Figure 24 depicts several additional generated images that exhibit semantic issues. For instance, one image features a person with his head and feet pointing in opposite directions, while another displays a fragmented pizza that does not cohere into a single entity. In yet another image, a blank painting hangs on the wall, creating a confusing and nonsensical composition.
@@ 3.5 Text, Noise, and Details
Noise, Color, and Blur Artifacts. Digital distortion in the form of pixelation or imperfect coloring can be present in generated images, particularly around the image edges. Monochrome areas may display semi-regular noise with horizontal or vertical banding, potentially due to the network attempting to replicate cloth textures. Older GANs tend to produce a more noticeable checkerboard noise pattern. Other telltale signs of deepfakes include inconsistencies in color or tone, oversaturation or undersaturation of colors, and unnatural image noise patterns. See the top row in Figure 26. Fluorescent bleed, where bright colors bleed onto the hair or face of a person in the image from the background, is also a potential indicator of a deepfake (the bottom row in Figure 26). The human attention system is naturally adept at quickly recognizing these patterns, making them useful tools for identifying generated images.
@@
Technical Details. AI-generated images may contain technical details that are either incorrect or appear as random shapes. For example, furniture legs can be particularly challenging for AI to accurately render, resulting in incorrect numbers of legs or physically impossible configurations. These issues can be attributed to the inherent difficulty of modeling complex objects and the limitations of the AI’s understanding of real-world. Some example failures are shown in Figure 28.
@@
Accurately rendering all details in complex scenes or crowd scenes, such as those depicted in Figures 29 and 30, can be particularly challenging for AI. The complexity of these scenes makes it difficult for the AI to accurately model every detail and can lead to errors in object placement, lighting, perspective, and other features. Despite the challenges, AI technology continues to improve, and advancements are being made in the generation of more realistic and believable large and crowd scenes.
@@ 4.1 Other Cues
In addition to the cues discussed above, there are several other indicators that can be used to identify deep fakes. One such method involves examining the metadata of an image or conducting a reverse Google search to verify its authenticity. Additionally, common sense can be applied to detect images that are likely to be generated, such as a shark swimming down a street or aliens eating sushi in a Chinese restaurant. Other indications of a deep fake include lack of motion blur, unnatural bokeh, all objects appearing in focus, and repeated patterns in the image.

### v5

# In what ways do text to image models fail?

@@ Abstract
The remarkable advancement of image and video generation models has led to the creation of exceptionally realistic content, posing challenges in differentiating between genuine and fabricated instances in numerous scenarios. However, despite this progress, a gap remains between the quality of generated images and those found in the real world. To address this, we have reviewed a vast body of literature from both academic publications and social media to identify qualitative shortcomings in image generation models, which we have classified into five categories. By understanding these failures, we can identify areas where these models need improvement, as well as develop strategies for detecting generated images and deepfakes. The prevalence of deepfakes in today’s society is a serious concern, and our findings can help mitigate their negative impact. In order to support research in this field, a collection of instances where models have failed is made available at here.
@@ 1 Introduction
Generated images, also known as synthetic images, are created by machine learning algorithms or other software programs, while real images are captured by cameras or other imaging devices. Generated images are not real-world representations of a scene or object, but rather computer-generated approximations. As such, they lack the authenticity and realism of real images. Deepfakes refer to fabricated media content that has undergone digital alterations to effectively substitute the appearance of one individual with that of another, creating a highly convincing outcome. This paper investigates the indicators that can be utilized for identifying artificially generated images, with a specific focus on detecting deepfakes.
@@
As the quality of generated images continues to improve, it is crucial to conduct more in-depth and precise analyses. Thus far, people have been amazed by the ability of synthesized images to approximate natural scenes. When Photoshop was introduced, significant efforts were made to identify manipulated images, and a similar approach is needed for generated images today. It would be beneficial to compile a set of indicators and other resources to aid in detecting generated images and deepfakes.
We present a collection of indicators that can be examined in a single image to determine whether it is genuine or generated. Overall, we offer five classes of these indicators including Human and Animal Body Parts, Geometry, Physics, Semantics and Logic, as well as Text, Noise, and Details, for both portraits and natural landscapes. The advantage of utilizing qualitative cues is that they are easily accessible and can be utilized by anyone, potentially serving as the initial step in detecting deepfakes.
@@ 2 Related work
2.1 Quantitative and Qualitative Approaches to Evaluate Generative Models
@@
Quantitative approaches have emerged as a vital tool to evaluate the performance of generative models. These methods rely on quantitative measures to assess how well a model is able to generate realistic data. One commonly used metric is the Inception Score [33], which evaluates the diversity and quality of generated images based on the classification accuracy of a pre-trained classifier. Another popular approach is the Fréchet Inception Distance [17], which uses feature statistics to compare the distribution of generated data with that of real data. Moreover, other metrics such as precision and recall [32] can be used to evaluate the quality of generated samples in specific domains such as vision, text and audio. Some studies have proposed methods to assess the visual realism of generated images (e.g.  [12]). These quantitative approaches provide a rigorous and objective way to measure the effectiveness of generative models, helping researchers to improve their models and develop more advanced generative techniques.
Recently, two metrics have gained popularity, namely the CLIP score and the CLIP directional similarity (e.g.  [27, 28]). The CLIP score evaluates the coherence of image and caption pairs by measuring their compatibility. A higher CLIP score indicates a greater degree of compatibility, which can also be interpreted as the semantic similarity between the image and the caption. Moreover, studies have shown that the CLIP score has a strong correlation with human judgement. On the other hand, the CLIP directional similarity is used for generating images based on text prompts while being conditioned on an input image. It assesses the consistency between the differences in the two images (in CLIP space) and the differences in their respective captions.
To obtain a thorough analysis of quantitative metrics for evaluating generative models, please refer to the following references [4, 6, 35, 38].
Qualitative assessment of generated images entails a human evaluation. The quality of these images is evaluated on various criteria, such as compositionality, image-text alignment, and spatial relations. DrawBench and PartiPrompts are prompt datasets used for qualitative benchmarking, that are were introduced by Imagen [31] and Parti [37], respectively. These benchmarks allow for side-by-side human evaluation of different image generation models.
PartiPrompts is a rich set of over 1600 prompts in English. It can be used to measure model capabilities across various categories and challenge aspects such as “Basic”, “Complex”, “Writing & Symbols”, etc.
Category Description Examples
Colors Ability to generate objects “A blue colored dog.”
with specified colors. “A black apple and a green backpack.”
Counting Ability to generate specified “Three cats and one dog sitting on the grass.”
number of objects. “Five cars on the street.”
Conflicting Ability to generate conflicting “A horse riding an astronaut.”
interactions b/w objects. “A panda making latte art.”
DALL-E [29] Subset of challenging prompts “A triangular purple flower pot.”
from [29]. “A cross-section view of a brain.”
Description Ability to understand complex and long “A small vessel propelled on water by oars, sails, or an engine.”
text prompts describing objects. “A mechanical or electrical device for measuring time.”
Marcus et al. [23] Set of challenging prompts “A pear cut into seven pieces arranged in a ring.”
from [23]. “Paying for a quarter-sized pizza with a pizza-sized quarter.”
Misspellings Ability to understand “Rbefraigerator.”
misspelled prompts. “Tcennis rpacket.”
Positional Ability to generate objects with “A car on the left of a bus.”
specified spatial positioning. “A stop sign on the right of a refrigerator.”
Rare Words Ability to understand rare words [^1] https://www.merriam-webster.com/topics/obscure-words. “Artophagous.”
“Octothorpe.”
Reddit Set of challenging prompts from “A yellow and black bus cruising through the rainforest.”
DALLE-2 Reddit [^2] https://www.reddit.com/r/dalle2/. “A medieval painting of the wifi not working.”
Text Ability to generate quoted text. “A storefront with ’Deep Learning’ written on it.”
“A sign that says ’Text to Image’.”
Table 1: Description and examples of the 11 categories in DrawBench, compiled from [31].
DrawBench is comprised of a collection of 200 prompts that are divided into 11 categories (Table 1), which aim to assess various capabilities of models. These prompts test a model’s ability to accurately render different attributes, such as colors, object counts, spatial relationships, text in the scene, and unusual object interactions. Additionally, the categories include complex prompts that incorporate lengthy, intricate textual descriptions, as well as uncommon words and misspelled prompts. DrawBench was used to directly compare different models, where human evaluators were presented with two sets of images, each consisting of eight samples, one from Model A and the other from Model B. Evaluators were then asked to compare Model A and Model B based on sample fidelity and image-text alignment.
Large-scale datasets have also been used in studies that focus on the qualitative evaluation of generated images (e.g.  [2]).
The assessment of models through qualitative methods can be susceptible to errors, potentially leading to an incorrect decision. Conversely, quantitative metrics may not always align with image quality. Therefore, the use of both qualitative and quantitative evaluations is typically recommended to obtain a more robust indication when selecting one model over another.
@@ 2.2 Deepfake Detection Methods
Detection of deepfakes has become an essential area of research due to the increasing sophistication of deep learning algorithms that can generate highly realistic fake images, videos, and audio. As a result, numerous deepfake detection methods have been proposed in recent years, ranging from traditional image and video forensic techniques to advanced deep learning-based approaches. These methods can be broadly categorized into two groups: static and dynamic analysis.
@@ 3 Qualitative Failures of Image Generation Models
We compiled a list of qualitative failures by examining images from various sources including social media websites such as Twitter, LinkedIn, Discord, and Reddit [^3] A few of the images used in this work were obtained with the consent of a Reddit user named Kronzky., as well as images from the DiffusionDB dataset [36] [^4] This dataset includes prompts that were used to generate images.. These images have been generated by notable generative models such as DALL-E 2, Midjourney, StableDiffusion, and Bing Image Creator. Additionally, we analyzed images from websites such as thisxdoesnotexist.com, whichfaceisreal.com, the Adobe Stock library, and openart.ai. We made sure that the text prompts used to generate images were not intentionally seeking peculiar images. Finally, we manually reviewed the images and filtered out the ones without problems.
@@ 3.1 Human and Animal Body Parts
Image Background. When creating generated images and deepfakes, issues with the background of the images may arise, particularly in cases where the face is in focus while the surrounding clues are incorrect. The neural network used to generate the images focuses mainly on the face and may not pay as much attention to the surrounding details. This can lead to strange companions or chaotic forms in the background. Additionally, the objects or people next to the primary person in the image may appear unnatural or “mutant". Figure 3 displays several instances of failures as examples.
@@
Figure 4: Here are some instances of eyes that were generated poorly. The eye in the bottom right corner is an actual photograph of a patient who has an irregularly shaped pupil. You can refer to this link for more details. This case represents a unique manifestation of a condition known as “cat’s eye Adie-like pupil," which is considered a warning sign for ICE syndrome.
@@ 3.2 Geometry
Generated images may exhibit anomalous or atypical image geometry, with objects appearing to be of an unusual shape or size, in comparison to their expected proportions.
Straight Lines and Edges. AI-generated images may lack the straight lines, seams, and connections found in real-world objects, resulting in wavy, misaligned, and jumpy renderings (e.g. in tiles). Generated images can also exhibit inconsistent or unnatural image edges, which refer to the boundaries between different parts of the image. Further, surfaces, which are typically straight, may look somewhat uneven in generated images. Some samples failures are shown in Figure 14.
@@
Perspective. Models lack the ability to understand the 3D world, which results in physically impossible situations when objects cross different planes in a scene. These errors are difficult to detect as our brain often auto-corrects them, requiring a conscious investigation of each angle of the object to identify inconsistencies. Generated images can display an unnatural or distorted perspective, where a person’s body appears stretched or compressed unrealistically. They may also have inconsistent or unrealistic camera angles, where a person’s face appears to be viewed from an impossible angle or perspective. Some example failures are shown in Figure 15.
@@
Other Geometry. Generated images exhibit various geometrical anomalies that may reveal their artificiality. For instance, their depth cues can be inconsistent or unnatural, causing the foreground or background to seem blurry or devoid of detail. Moreover, they often lack parallax, which is the apparent displacement of objects when viewed from different perspectives, resulting in a flat or two-dimensional appearance. Additionally, incorrect or inconsistent motion blur may suggest that certain parts of the image have been manipulated. The absence of occlusion, i.e., the overlapping of objects in the scene, is another telltale sign of generated images, as it can make the image look flat or unrealistic. Lastly, generated images may display improper image alignment, with objects seeming misaligned or out of place.
@@ 3.3 Physics
Reflection. An effective technique for detecting generated images is to examine the lighting and how it interacts with the elements within the image, and how it causes reflections and shadows. Generated images can exhibit artificial reflections that are inconsistent with the natural lighting and environment, such as those in glasses, mirrors, or pupils. The root cause of this issue is that deep generative models lack a proper understanding of reflections. While these models may recognize that an image contains a reflection and typically involves two people (one facing the camera and the other with their back turned), they do not comprehend that the two individuals are, in fact, the same person. Generated images may display other lighting effects that do not match real-world environments, such as lens flares, lens distortion, chromatic aberration and unnatural specular highlights. These effects are frequently observed in genuine photographs due to the physical properties of camera lenses and the way light is refracted through them. Some example failures are shown in Figure 19.
@@
AI-generated images sometimes exhibit inconsistencies in the geometry of reflections [13]. Shown in the left panel of Fig. 20 is a photographic image in which lines that connect points on the toy dinosaur in the scene and their reflections in the mirror all converge to a single point (a vanishing point). Reflections in AI-generated images (the right panel in Fig. 20), however, exhibit physical inconsistencies as can be seen by the lack of a consistent vanishing point.
Figure 20: Consistent and inconsistent reflections in real (left) vs. generated images.
Shadow. Generated images might not include shadows, which are typically found in real images and contribute to the impression of depth and authenticity. It is important to observe objects without shadows and those with highlights that appear to originate from a different direction than the rest of the image. Additionally, if the photo was taken outdoors in natural light during the afternoon, the setting sun will produce longer shadows than it would at midday, which can be easily identified by scrutinizing the shadow’s length. However, this method may not be as precise in artificial lighting conditions. Finally, if there are multiple objects or people within the scene, their shadows should be consistent with each other. Some generated images with inconsistent shadows are shown in 21.
@@ 3.4 Semantics and Logic
Figure 23: Samples of spatial reasoning from [23]. Images are generated by DALL-E 2 for the following text prompts for columns from left to right: “a red basketball with flowers on it, in front of blue one with a similar pattern", “a red ball on top of a blue pyramid with the pyramid behind a car that is above a toaster", “a pear cut into seven pieces arranged in a ring, “In late afternoon in January in New England, a man stands in the shadow of a maple tree", and “An old man is talking to his parents".
Context and Scene Composition. Generated images can be detected through various inconsistencies such as the background or surroundings not matching the real-world environment, cardinality/counting, missing contextual details, unnatural object placement, and inconsistent image composition. These irregularities may include inconsistencies in order of objects, missing objects or features, objects appearing in the wrong location or orientation, or unnatural arrangement and placement of objects in the image. Please see Figure 24.
@@
Other Semantics. Figure 25 depicts several additional generated images that exhibit semantic issues. For instance, one image features a person with his head and feet pointing in opposite directions, while another displays a fragmented pizza that does not cohere into a single entity. In yet another image, a blank painting hangs on the wall, creating a confusing and nonsensical composition. From time to time, models encounter issues when creating reverse images. At instances, these models produce highly similar objects or faces in the images (refer to Figure 26).
@@
Figure 26: Further images with semantic problems, top: problem with generating inverse images, bottom: sometimes models generate similar objects of faces in the images.
@@ 3.5 Text, Noise, and Details
Noise, Color, and Blur Artifacts. Digital distortion in the form of pixelation or imperfect coloring can be present in generated images, particularly around the image edges. Monochrome areas may display semi-regular noise with horizontal or vertical banding, potentially due to the network attempting to replicate cloth textures. Older GANs tend to produce a more noticeable checkerboard noise pattern. Other telltale signs of generated images include inconsistencies in color or tone, oversaturation or undersaturation of colors, and unnatural image noise patterns. See the top row in Figure 28. Fluorescent bleed, where bright colors bleed onto the hair or face of a person in the image from the background, is also a potential indicator of a generated images (the bottom row in Figure 28). The human attention system is naturally adept at quickly recognizing these patterns, making them useful tools for identifying generated images.
@@
Fine-grained Details. AI-generated images may contain technical details that are either incorrect or appear as random shapes. For example, furniture legs can be particularly challenging for AI to accurately render, resulting in incorrect numbers of legs or physically impossible configurations. These issues can be attributed to the inherent difficulty of modeling complex objects and the limitations of the AI’s understanding of real-world. Some example failures are shown in Figure 30.
@@
Accurately rendering all details in complex scenes or crowd scenes, such as those depicted in Figures 31 and 32, can be particularly challenging for AI. The complexity of these scenes makes it difficult for the AI to accurately model every detail and can lead to errors in object placement, lighting, perspective, and other features. Despite the challenges, AI technology continues to improve, and advancements are being made in the generation of more realistic and believable large and crowd scenes.
@@ 4.1 Other Cues
In addition to the cues discussed above, there are several other indicators that can be used to identify generated images and deepfakes. One such method involves examining the metadata of an image or conducting a reverse Google search to verify its authenticity. Additionally, common sense can be applied to detect images that are likely to be generated, such as a shark swimming down a street or aliens eating sushi in a Chinese restaurant. Other indications of generated images and deepfakes include lack of motion blur, unnatural bokeh, all objects appearing in focus, and repeated patterns in the image.